You are testing the null hypothesis that there is no relationship between two variables, X and Y. From your sample of n = 22, you determine that SSR = 80 and SSE = 20. Construct the ANOVA table, test for model significance at a 0.05 level of significance and state your conclusion, and calculate the coefficient of determination and interpret its meaning.
Answer #1

Since there are two are variables so degree of freedom of regression is

The degree of freedom of error is

Following is the completed ANOVA:

ANOVA
df SS MS F p-value
Regression 1 80 80/1=80 80/1=80 0
Error 20 20 20/20=1
Total 21 100

For p-value excel function used is : =FDIST(80,1,20)

Since p-value is less than 0.05 so model is significant.

----------------------------------------------

The R-square is

It shows that 80% variation in dependent variable is explained by independent variable.